Grade 11Physical Chemistry

How many litres of liquid ccl4 (d=1.5g/cc) must be measured out to contain 1*10^25 cl atoms??

Dev Tandon
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago

In ccl4  four molecules of cl are present,
So 1* 10^25 atoms = ¼ * 1*10^25 = 25 x 1023 molecules .
Mass of 6.023 molecules of chlorine = 153.8 ( molar mass of ccl4)
So mass of 25 x 10 3 molecules = 153.8 / 25 x 10 23 = 638.57g
Density of ccl4 = 1.5g/cc
Volume of ccl4 to be measured = 638.57/1.5 = 425.17 cc